OpenStudy (technodynamic):

Let f(x)=x^2+1 and h(x)=sqrt(2x) . Find a formula for f(h(x))=

OpenStudy (anonymous):

consider sqrt(2x) as your x in f(x), so f(h(x)) should be f(sqrt(2x))=sqrt(2x)^2+1=2x+1

OpenStudy (kira_yamato):

\[f(h(x)) = [\sqrt{2x}]^2 + 1 = 2x +1\]

OpenStudy (anonymous):

Yours is better.